Suppose a country decreases income taxes by $300 billion, and this leads to an increase in consumption spending of $150 billion. Suppose the multiplier is 1.5 and the economy’s real GDP is $5,000 billion.

  • In which direction will the aggregate demand curve shift and by how much?
Answer #1

Taxes are reduced by 300 billion. Consumption is increased by 150 million. This implies that disposable income is increased by 300 billion and so MPC = 150/300 = 0.5. Now tax multiplier is -MPC/(1-MPC) = -0.5/(1-0.5) = -1. Hence, when taxes are reduced by 300 billion, income or real GDP will also increase by 300 x 1 = 300 billion.  Aggregate demand curve shift to the right when taxes are reduced and the size of shift is 300 billion.